OpenStudy (anonymous):

There are 6 girls and 5 boys in a math class. If the teacher chooses two students at random to answer questions, what is the probability of selecting a boy then a girl?

OpenStudy (anonymous):

I really need help in this one! possible answers: 3/11 36/121 6/11

OpenStudy (anonymous):

Total Students in the class: 6 + 5 = 11 Now, firstly, find the probability of selecting a boy: P(B) = 5/11 Now, One boy is selected, so number of students left behind = 11 - 1 = 10 So, from total of 10 students, Find the probability of choosing a Girl: P(G) = 6/10 So, Probability of Selecting a boy and then a girl will be: \[P(A)*P(B) = \frac{5 \times 6}{11 \times 10}\] \[= \frac{30}{110} = \frac{3}{11}\]

OpenStudy (anonymous):

what will be the probability of selecting a boy and then a girl?

OpenStudy (anonymous):

3/11

OpenStudy (anonymous):

thanks!! you saved me!

OpenStudy (anonymous):

Welcome dear...

OpenStudy (anonymous):

can I ask you one more small question?

OpenStudy (anonymous):

Sure...

OpenStudy (anonymous):

Why are you asking a question before asking a question.. Ha ha ha ha...

OpenStudy (anonymous):

Which of the following simplifies to

OpenStudy (anonymous):

jajajaj funny!

OpenStudy (anonymous):

I could not get your question.. Following are what???

OpenStudy (anonymous):

I don't get in neither :(

OpenStudy (anonymous):

I have to simplify that

OpenStudy (anonymous):

the possible answers are

OpenStudy (anonymous):

and

OpenStudy (anonymous):

See,